OpenStudy (anonymous):

Brenda is training to be a volunteer at her library. She wonders what the probability is that the first item she'll get to check out for a customer will be a video. So, on her first day of training, she tracks what items are checked out. The results are shown here in a tally chart. Books:37 Videos:24 Magazines:19 Use Brenda's experimental data to estimate the probability that her first item to check out will be a video. Show and explain your work.

OpenStudy (anonymous):

add up everything get 37+24+19 = 80

OpenStudy (anonymous):

there are 24 videos, and the ratio of videos to total is \[\frac{24}{80}=\frac{3}{10}=.3\]